# Percolation analysis of spatio-temporal distribution of population in Seoul and Helsinki

### v1

@@ IV Discussion
In summary, we identify the macroscopic structure underlying the spatial distribution of near-real time population in Seoul and Helsinki with a percolation framework. We observe that the growth patterns of the largest clusters in the real population data and their randomized data are significantly different, and highly populated areas during daytime are denser and form larger clusters than during nighttime. We investigated the cluster size distributions at percolation criticality and observed that their power-law exponents during daytime are lower than during nighttime, indicating that spatial distributions of urban population during daytime and nighttime fall into different universality class. We measured the area-perimeter fractal dimension of the collection of clusters at percolation criticality and observed that the fractal dimensions during daytime are higher than during nighttime, indicating that the perimeters of high-population clusters during daytime are more rough than during nighttime. Our findings suggest that even the same city can have qualitatively different spatial distributions of population over time, and propose a way to quantitatively compare the macrostructure of cities based on population distribution data.
What makes such qualitative differences between daytime and nighttime population distributions in the same city? A possible answer to this question is commuting as people concentrate in a few small areas (e.g., urban centers) during daytime and disperse after work. However, more quantitative analysis based on commuting data is necessary to check the validity of this hypothesis and this will be one of immediate future works of our study. Comparing our results from mobile phone users to the results from the registered population collected by census will be also interesting as nighttime population distribution is often assumed to be similar to the registered population distribution.
We argue that our framework minimizes the influence of microscopic details of the system such as grid cell size since it focuses on the critical exponent and the fractal dimension at criticality. Our percolation-based analysis of mascrostructure of dynamic urban population can provide an informative method for quantitative understanding of where people are in cities, how we can compare the spatial distributions of population between cities or between times.

### v2

@@ IV Discussion
In summary, we identified the macroscopic structure underlying the spatial distribution of near-real time population in Seoul and Helsinki with a percolation framework. We observe that the growth patterns of the largest clusters in the real population data and their randomized data are significantly different, and highly populated areas during daytime are denser and form larger clusters than during nighttime. We investigated the cluster size distributions at percolation criticality and observed that their power-law exponents during the daytime are lower than during the nighttime, indicating that spatial distributions of urban population during daytime and nighttime fall into different universality class. We measured the area-perimeter fractal dimension of the collection of clusters and observed that the fractal dimensions during daytime are higher than during nighttime, indicating that the perimeters of high-population clusters during daytime are more rough than during nighttime. Our findings suggest that even the same city can have qualitatively different spatial distributions of population over time, and propose a way to quantitatively compare the macrostructure of cities based on population distribution data.
What makes such qualitative differences between daytime and nighttime population distributions in the same city? A possible answer to this question is commuting as people concentrate in a few small areas (e.g., urban centers) during daytime and disperse after work. However, more quantitative analysis based on commuting data is necessary to check the validity of this hypothesis and this will be one of immediate future works of our study. Mathematical models such as fractional Brownian motion models may help address this question as in the other real-world application of percolation [41]. Active-particle models can also be a good candidate for future research on the emergence of the observed patterns in the spatial distributions of urban population. Finally, one can investigate population distributions in other cities with different geometries, geographies, and socio-economic conditions with our framework to check whether there is a universal pattern across cities or whether cities can be classified into different groups [20]. Comparing our results from mobile phone users to the results from the registered population collected by census will be also interesting as nighttime population distribution is often assumed to be similar to the registered population distribution.
We argue that our framework minimizes the influence of microscopic details of the system such as grid cell size since it focuses on the critical exponent and the fractal dimension. But our framework assumes that population density is spatially heterogeneous and high-density areas are clustered in cities. In addition, if the population of a city is too sparely distributed in space, population clusters are not well defined. Our percolation-based analysis of macrostructure of dynamic urban population can provide an informative way to quantitatively understand the location of populations and compare the spatial distribution of population across cities or over time. Our findings suggest that urban planners and policy makers need to consider the differences in demand between daytime and nighttime when allocating resources and infrastructure in cities.